# WAL-Net: Weakly supervised auxiliary task learning network for carotid plaques classification

### 25 Jan 2024 v1

@@ 4 Experiments and Results
Fig. 3: The preprocessing pipeline for carotid artery plaque ultrasound images. (a): The original ultrasound images of the three types of plaques: hyperechoic plaque, hypoechoic plaque, and mixed-echoic plaque. (b): The regions of interest in the ultrasound images of the three types of plaques. (c): The ultrasound images of the three types of plaques after being uniformly resized.
The data used for evaluating experiments in this paper is derived from a carotid artery ultrasound image dataset obtained from Zhongnan Hospital of Wuhan University. The study received approval from the Institutional Review Board (IRB) of the hospital. Each lesion image in the dataset is accompanied by a corresponding lesion category label. The dataset comprises a total of 1,270 carotid artery ultrasound images collected from 844 patients by ultrasound imaging experts. Among these images, there are 301 hyperechoic plaque images, 605 hypoechoic plaque images, and 364 mixed-echoic plaque images. For evaluation and experimentation purposes, the dataset is split into training, validation, and test sets in a 6:2:2 ratio.
@@ 4.2 Experimental Setup
The experiments conducted in this study were implemented using the Python programming language and the PyTorch framework. The optimizer used in the experiments was the Adam optimizer with a learning rate of 0.0001. A batch size of 8 was selected for the experiments. All experimental results are the averages obtained after five random experiments on the same device. Five evaluation metrics were defined for the classification of carotid artery plaque ultrasound images, including accuracy, F1-score, kappa, precision, and recall.
@@ 4.3 Comparison with Other Classification Methods
Sequencer (2022)  (tatsunami2022sequencer, )
@@
Fig. 5: Confusion matrix for WAL-Net and baseline.
@@ 4.4 Visualization of Weakly Supervised Segmentation Results
Fig. 6: Visualization examples of the generated pseudo mask and the predicted results by proposed WAL-Net.
@@ 4.6 Comparison with Different ROI Augmentation Methods
In the work conducted by Amirreza Mahbod et al.  (mahbod2020effects, ), various methods influencing segmentation for classification were experimentally explored, and a particularly effective non-end-to-end approach was identified. Extending this methodology to end-to-end networks, as demonstrated by He et al.  (he2023joint, ), involved similar operations on the advanced features of classification networks, yielding favorable outcomes. To validate the applicability of this approach to the specific context of carotid artery ultrasound image datasets, we conducted experiments employing different segmentation strategies. As shown in Table 3, ’bg rm’ represents the removal of background values from features, ’bg rm&crop’ represents eliminating background values and crop the foreground to a fixed size, ’crop’ signifies no removal of background values but direct crop of the foreground to a fixed size, and ’rwm’ refers to the method proposed by  (fu2023smc, ), which multiplies the segmentation output predictions with advanced features for classification, emphasizing foreground weights while diminishing background weights. ’dilated crop’ represents cropping the foreground and a portion of the surrounding background, then resizing to a fixed size, which is the method adopted in this paper. The results indicate the superiority of our proposed method over various strategies in the given context.
@@ 5 Conclusion
In this paper, we posit that judiciously harnessing the intrinsic correlations between different tasks in auxiliary task learning is crucial for improving the classification results of carotid plaque ultrasound images. Consequently, we introduce a novel Weakly Supervised Auxiliary Task Learning Network (WAL-Net) comprising a shared encoder, a classification task head, and a weakly supervised segmentation task head. In contrast to traditional classification approaches, WAL-Net incorporates an auxiliary task based on weakly supervised learning, specifically, the segmentation task. Exploiting the auxiliary task, we explicitly enhance the classification task using the RCM module, thereby improving the performance of the classification task. We also design a module for the supervision of weakly supervised auxiliary task learning, utilizing the combination of unsupervised learning and attention mechanisms to generate pseudo-segmentation labels, thereby completely alleviating dependence on real segmentation labels while still achieving satisfactory segmentation results. Various experiments on the carotid ultrasound dataset demonstrate the effectiveness of our approach. It is worth noting that optimizing the PGM module to obtain better pseudo-segmentation labels without relying on real labels will make the auxiliary task more effective and precise. In future work, we plan to expand the PGM module to enhance the performance of the auxiliary task. Furthermore, while our method is specifically applied to carotid plaque ultrasound images, WAL-Net and its individual sub-modules proposed herein are generalizable and hold potential for application in other image classification domains.

### 27 Jan 2024 v2

@@ 4 Experiments and Results
Figure 3: The preprocessing pipeline for carotid artery plaque ultrasound images. (a): The original ultrasound images of the three types of plaques: hyperechoic plaque, hypoechoic plaque, and mixed-echoic plaque. (b): The regions of interest in the ultrasound images of the three types of plaques. (c): The ultrasound images of the three types of plaques after being uniformly resized.
The data used for evaluating experiments in this paper is derived from a carotid artery ultrasound image dataset obtained from Zhongnan Hospital of Wuhan University. The study received approval from the Institutional Review Board (IRB) of the hospital. Each lesion image in the dataset is accompanied by a corresponding lesion category label. The dataset comprises a total of 1,270 carotid artery ultrasound images collected from 844 patients by ultrasound imaging experts. Among these images, there are 301 hyperechoic plaque images, 605 hypoechoic plaque images, and 364 mixed-echoic plaque images. For evaluation and experimentation purposes, the dataset is split into training, validation, and test sets in a 6:2:2 ratio. The code used in the experimental section of this paper has been uploaded to ’https://github.com/a610lab/WAL-Net’
@@ 4.2 Experimental Setup
The experiments conducted in this study were implemented using the Python programming language and the PyTorch framework. The optimizer used in the experiments was the Adam optimizer with a learning rate of 0.0001. A batch size of 8 was selected for the experiments. All experimental results are the averages obtained after five random experiments on the same device.
Five evaluation metrics were defined for the classification of carotid artery plaque ultrasound images, including accuracy, F1-score, kappa, precision, and recall.
@@ 4.3 Comparison with Other Classification Methods
Sequencer (2022)  (Tatsunami and Taki, 2022)
@@
Figure 5: Confusion matrix for WAL-Net and baseline.
@@ 4.4 Visualization of Weakly Supervised Segmentation Results
Figure 6: Visualization examples of the generated pseudo mask and the predicted results by proposed WAL-Net.
@@ 4.6 Comparison with Different ROI Augmentation Methods
In the work conducted by Amirreza Mahbod et al.  (Mahbod et al., 2020), various methods influencing segmentation for classification were experimentally explored, and a particularly effective non-end-to-end approach was identified. Extending this methodology to end-to-end networks, as demonstrated by He et al.  (He et al., 2023), involved similar operations on the advanced features of classification networks, yielding favorable outcomes. To validate the applicability of this approach to the specific context of carotid artery ultrasound image datasets, we conducted experiments employing different segmentation strategies. As shown in Table 3, ’bg rm’ represents the removal of background values from features, ’bg rm&crop’ represents eliminating background values and crop the foreground to a fixed size, ’crop’ signifies no removal of background values but direct crop of the foreground to a fixed size, and ’rwm’ refers to the method proposed by  (Fu et al., 2023), which multiplies the segmentation output predictions with advanced features for classification, emphasizing foreground weights while diminishing background weights. ’dilated crop’ represents cropping the foreground and a portion of the surrounding background, then resizing to a fixed size, which is the method adopted in this paper. The results indicate the superiority of our proposed method over various strategies in the given context.
@@ 5 Conclusion
In this paper, we posit that judiciously harnessing the intrinsic correlations between different tasks in auxiliary task learning is crucial for improving the classification results of carotid plaque ultrasound images. Consequently, we introduce a novel Weakly Supervised Auxiliary Task Learning Network (WAL-Net) comprising a shared encoder, a classification task head, and a weakly supervised segmentation task head. In contrast to traditional classification approaches, WAL-Net incorporates an auxiliary task based on weakly supervised learning, specifically, the segmentation task. Exploiting the auxiliary task, we explicitly enhance the classification task using the RCM module, thereby improving the performance of the classification task. We also design a module for the supervision of weakly supervised auxiliary task learning, utilizing the combination of unsupervised learning and attention mechanisms to generate pseudo-segmentation labels, thereby completely alleviating dependence on real segmentation labels while still achieving satisfactory segmentation results. Various experiments on the carotid ultrasound dataset demonstrate the effectiveness of our approach.
It is worth noting that optimizing the PGM module to obtain better pseudo-segmentation labels without relying on real labels will make the auxiliary task more effective and precise. In future work, we plan to expand the PGM module to enhance the performance of the auxiliary task. Furthermore, while our method is specifically applied to carotid plaque ultrasound images, WAL-Net and its individual sub-modules proposed herein are generalizable and hold potential for application in other image classification domains.